OpenStudy (anonymous):

Solve 3x2 + 5x + 7 = 0. Round solutions to the nearest hundredth. No real solutions x ≈ −6.34 and x ≈ 1.34 x ≈ −0.45 and x ≈ 2.11 x ≈ −2.11 and x ≈ 0.45

OpenStudy (anonymous):

roots of equation: \[ax^2+bx+c=0\] are \[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

now we have a=3 , b=5 and c=7 what will u get if u put them in the formula?

OpenStudy (anonymous):

notice \[b^2-4ac=5^2-4*3*7=25-84=-59\] is a negative number